MODS1 Commissioning Run 3: UTC 2010 December 19

Observer(s): R. Pogge and C. Henderson (OSU), M. Pedani (LBTO)
Telescope Operator: D. Gonzalez (LBTO)
Telescope Support: J. Little (LBTO)
Instrument Support: J. Morris (LBTO)

Summary

We were in 100% humidity with drippy conditions just above freezing all night. We haven't opened yet in 4 nights...

Worked on scripting and did some tests of light leaks with enclosure lights on and off while taking biases and direct images whose length it comparable to sets of calibration images (esp. flats). Also ran a dark sequence on both channels, and tested some of the start/stop/sleep/wake procedure (.pro) scripts.

Preparations

Minor modifications to the scripting interface to move some of the raw engineering output behind a "verbose" flag to get the feel closer to what an observer would expect to see from a running script.

Fixed some minor issues in the MODS user interface uncovered during the previous nights. Prep for testing acquisition sequences.

Details

All times are UTC unless otherwise noted.

Afternoon

We've been in and out of clouds - periods of clear showing very high cirrus bands, otherwise we're about level with surrounding clouds. Humidity has been 100% all day. Wind started to increase through the afternoon with gusts up to 20m/s by 3pm MST. Temperature has been hovering around 0°C since about noon.

If conditions improve enough to open up during the last half or so of the night the seeing is likely to be poor. As such, our main plan if we open is to work on scripting the target acquisition process (during past runs nearly all acquisitions were "by hand"). This includes tests of when in the preset sequence to lock/center the IMCS, rooting out bugs in the acquisition scripts, etc.

0020 - Clouds blowing over the summit at sunset, providing enough attenuation to look at the disk of the sun safely and watch as it disappears behind the distant mountains. Temperature still hovering around 0°C, mirror temp is +4.4°C, so if we get to open equilibration will be a problem.

0110 - Humidity 100% and right at 0°C. Winds around 15m/s from the SSW, gusts up to 20m/s. Cirrus overhead in bands with whisps of stuff blowing by at ground level every now and then.

0200 - Humidity and Temperature haven't budged a bit in the past hour (100% and 0°C), but the wind is not as strong, now 10-15m/s and out of the SW. High clouds all around the sky, and though we're no longer enshrouded in blowing clouds, there is condensation everywhere.

0320 - Conditions unchanged, humidity still pegged and temperatures hovering around 0°C. Mostly overcast with fast-moving high clouds.

0415 - Still cloudy, humidity stil 100% and slightly above freezing. Might be some relief off to the west based on satellite IR and water vapor maps. Good thing winter nights are so long (also a bad thing winter nights are so long).

0500 - Enshrouded in wet, drippy fog, 100% humidity and still ~0°C. Called for software assistance as we started noticing that the local network is very slow to do things like launch new xterms, ssh between internal machines, etc.

0600 - Still enshrouded, just above freezing making the building drippy. Folks in Tucson investigating our network's case of the slows...

0700 - No change, except to be wetter and drippier out on the patio. The network is responding normally now, thanks!

0800 - Much the same, still sitting in a cloud: wind is 10-15m/s out of the SW, just above freezing and 100% humidity. It is very drippy as the air is very saturated, setting off the rain sensor.

0900 - Still in a drippy cloud. Temperature slowly falling towards freezing the past hour.

0912 - Started a long (~2h) dark sequence on both cameras, room lights off.

1000 - Wind picking up, gusts up to 22m/s in the past half hour, still from the SW. Humidity 100% and temps hovering around freezing. Still enshrouded in fast-moving cloud with occasional glimpses of the moon.

1100 - Wind now gusting up to 24m/s, humidity still pegged out, temp unchanged. Dark sequence in Red and Blue complete, including matching bias images.

1220 - Weather unchanged. Last bits done, MODS put to sleep for the day using the modsSleep.pro script. Will try waking with modsWake.pro later today.

-- RichardPogge - 18 Dec 2010
Topic revision: r7 - 19 Dec 2010, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback